OpenStudy (anonymous):

A movie rental store sells a couples movie pack for and a family movie pack for . In one year, the Nguyen family paid for 13 movie packs. How many family movie packs did they pay for?

OpenStudy (anonymous):

UGLYYYYYYYYYYYYYYY

OpenStudy (anonymous):

well now the question

OpenStudy (anonymous):

is this question complete .... i think you missed some info

OpenStudy (hba):

dollar 6 dollar 10 and dollar 126

OpenStudy (hba):

6x + 10y = 126.........(1) x + y = 13................(2) (2) into (1) gives: 6x + 10(13 - x) = 126 => 6x + 130 - 10x = 126 i.e. 4x = 4 => x = 1 => y = 12 so, the Nguyen's bought 12 family movie packs.

OpenStudy (anonymous):

rattay huye hain question tjhe

OpenStudy (hba):

kia karee

OpenStudy (anonymous):

ok these are what u can chose from 9 10 11 12

OpenStudy (hba):

12 hai na answer

OpenStudy (hba):

bewakoof

OpenStudy (anonymous):

@barbieodog : you are ugly and you dont have brains .... 12 is the answer